40m
QIL
Cryo_Lab
CTN
SUS_Lab
TCS_Lab
OMC_Lab
CRIME_Lab
FEA
ENG_Labs
OptContFac
Mariner
WBEEShop
|
40m Log |
Not logged in |
 |
|
Tue Oct 8 18:35:09 2019, gautam, Update, LSC, Locking prep
|
Wed Oct 9 18:15:26 2019, gautam, Update, LSC, PRMI 3f locking
|
Wed Oct 9 22:02:58 2019, gautam, Update, LSC, REFL55 whitening issue
|
Thu Oct 10 01:12:56 2019, gautam, Update, LSC, Locking studies
|
Wed Oct 16 11:42:17 2019, gautam, Update, LSC, Poor separation of PRCL/MICH in 3f signals
|
Thu Oct 17 11:19:28 2019, gautam, Update, LSC, Locking activity last night 
|
Tue Oct 22 00:52:27 2019, gautam, Update, LSC, Locking updates
|
Tue Oct 29 01:24:45 2019, gautam, Update, LSC, More locking updates 
|
Tue Oct 29 15:13:19 2019, gautam, Update, LSC, More locking updates
|
Wed Oct 30 01:27:00 2019, gautam, Update, LSC, More locking updates
|
Wed Oct 30 11:53:41 2019, gautam, Update, LSC, MICH loop shape tuning 
|
Wed Oct 30 17:08:40 2019, gautam, Update, LSC, POP22 investigation
|
Mon Nov 4 15:29:47 2019, gautam, Update, LSC, POP signal path
|
Mon Nov 4 16:06:58 2019, gautam, Update, LSC, POP optical path    
|
Tue Nov 5 11:52:27 2019, gautam, Update, LSC, Locking notes
|
Wed Nov 6 02:08:48 2019, gautam, Update, LSC, Locking updates
|
Wed Nov 6 17:05:45 2019, gautam, Update, LSC, CARM calibration  
|
Wed Nov 6 17:45:34 2019, gautam, Update, LSC, ~
|
Mon Nov 18 21:04:38 2019, gautam, Update, LSC, Locking - some ideas
|
Wed Nov 20 01:07:18 2019, gautam, Update, LSC, Locking - progress
|
Wed Nov 20 12:14:17 2019, gautam, Update, LSC, Locking - progress
|
Wed Nov 20 17:20:24 2019, Yehonathan, Update, LSC, QPD Investigation
|
Wed Nov 20 17:52:00 2019, gautam, Update, LSC, QPD MEDM screen update  
|
Wed Nov 20 21:29:28 2019, gautam, Update, LSC, PRG ~13
|
Fri Nov 22 00:54:14 2019, gautam, Update, LSC, locking notes
|
Mon Nov 25 19:11:22 2019, gautam, Update, LSC, ALS noise re-look
|
Wed Nov 27 23:24:01 2019, gautam, Update, LSC, No shaking but no inspiration either
|
Sun Dec 1 13:38:48 2019, rana, Update, LSC, No shaking but no inspiration either
|
Tue Dec 3 00:03:57 2019, gautam, Update, LSC, Green ALS also shows elevated noise with high arm buildup
|
Wed Dec 4 19:54:27 2019, gautam, Update, LSC, A look to the past
|
Thu Dec 5 01:54:39 2019, gautam, Update, LSC, Partial CM board path engaged
|
Fri Dec 6 00:02:48 2019, gautam, Update, LSC, What is the correct way to set the 3f offsets?
|
Fri Nov 15 12:08:04 2019, gautam, Update, LSC, POPDC whitening board
|
Tue Oct 29 17:40:48 2019, gautam, Update, LSC, More locking updates 
|
|
Message ID: 15012
Entry time: Tue Nov 5 11:52:27 2019
In reply to: 15010
Reply to this: 15014
|
Author: |
gautam |
Type: |
Update |
Category: |
LSC |
Subject: |
Locking notes |
|
|
Summary:
I am still unable to achieve arm powers greater than TRX/TRY ~10 while keeping the PRMI locked. A couple of times, I was able to get TRY ~50, but TRX stayed at ~10, or even dropped a little, suggestive of a DARM offset? On the positive side, the ALS system seems to work pretty reliably, and I can keep the arms controlled by ALS for several tens of minutes.
Details:
- Despite my POP beam path improvements, I saw the POP22 level drop as I lowered the CARM offset.
- One strange feature last night was that with the arms held off resonance using ALS, I had to flip the sign and increase the gain by ~x2 of the REFL33_I-->PRCL loop in order to lock the PRMI. This was confirmed by locking on the 1f error signals and measuring the ratio of the response between the 1f and 3f signals while shaking PRCL using DTT swept sine.
- At different CARM offsets, I noted that the DC offset level on the 1f photodiodes (i.e. REFL11 and AS55) were changing significantly.
- I ran a measurement of the sensing matrix with the arm powers hovering around ~10, which is just before I lose the PRMI lock - managed to stay locked for >5 minutes, but the sensing matrix seems to suggest that the REFL33 demod angle needs to be rotated - maybe this is the reason why the PDH horn-to-horn voltage of REFL33 is lower now than it was last week? No idea why that should be, I was around the LSC rack but if the situation is so fragile, seems hopeless.
- MICH sensed by REFL165_Q still seems stable, so that's good...
- So my best hypothesis at the moment is that the PRCL optical gain is falling as I reduce the CARM offset (due to DC offset? or something else?). Needs some detailed modeling for more insight, I'm out of ideas for tests to run while locking as I've gone through the full gamut of OLTF and sensing matrix measurements at various CARM offsets without getting any clues as to what's going on.
|
|